# DisC-GS: Discontinuity-aware Gaussian Splatting

### v1

@@ 2 Related Work
Novel View Synthesis. Owing to the wide range of applications, the task of novel view synthesis has received lots of research attention [20, 37, 36, 39, 21, 33, 2, 3, 4, 40, 46, 6, 10, 15, 16, 34, 27, 23, 12, 45, 17, 42, 19, 32, 30, 29, 44, 48]. In the early days, with the emergence of CNN, different works have been proposed to leverage CNN in this task from different perspectives. Among them, Hedman et al. [20] proposed to use CNN to predict blending weights, and Sitzmann et al. [37] proposed to seek help from CNN in performing volumetric ray-marching. As time passed, NeRF tends to become a popular way in representing 3D scenes. Specifically, the original version of NeRF is first proposed by Mildenhall et al. in [33] and after it comes out, a variety of different NeRF-based methods have been further proposed, such as Mip-NeRF [2], NeRF++ [46], and Point-NeRF [40]. Despite the increased efforts, a weakness of NeRF-based methods can be that, to render novel-view images in high visual quality, they often still require a slow rendering process [27, 12]. This can negatively affect the usage of these methods in many real-world scenarios.
@@ 3 Preliminary
where $N$ represents the number of projected Gaussians that overlap $p$, $c_{i}$ represents the color of the $i$-th Gaussian calculated from its corresponding SH coefficients, $\alpha_{i}$ represents the opacity of the $i$-th Gaussian, $\mu^{2D}_{i}$ represents the center of the $i$-th projected Gaussian, and $\Sigma^{2D}_{i}$ represents the covariance matrix of the $i$-th projected Gaussian. Note that, no matter whether Gaussian Splatting represents the 3D scene using 3D or 2D Gaussian distributions, the above equations can describe its rendering process consistently. In fact, as also mentioned in [23], the difference between rendering images from 3D or 2D Gaussians can be reduced to that, when the scene is represented through 2D Gaussians, the scale matrix $S$ of each of the 2D Gaussians should contain a zero column vector. In this work, we apply our framework to both 2D and 3D Gaussian Splattings, achieving performance improvements as shown in Tab. 2. Yet, as pointed out by [23], using 3D Gaussians instead of 2D Gaussians to represent the scene can result in a viewpoint inconsistency problem. Thus, in Sec. 4, we focus on explaining how our framework is applied to 2D Gaussian Splatting, in which we fix the last column of the scale matrix $S$ of all the Gaussians to be a zero vector. We then discuss the application of our framework on 3D Gaussian Splatting in Sec. 5.2.
@@ 4 Proposed Method: DisC-GS
Yet, we here argue that Gaussian Splatting has a key intrinsic limitation: it may fail to render discontinuities and boundaries accurately. To tackle this problem, in this work, we propose a novel framework named DisC-GS, which can seamlessly equip Gaussian Spatting with the discontinuity rendering ability. Specifically, during rendering images from the 3D scene, to render discontinuities properly, DisC-GS enables each Gaussian distribution projected onto the image plane to be first “pre-scissored” along certain desired boundaries before being rendered. However, such a “pre-scissoring” operation by itself can break the differentiability of the framework. Considering this, we further incorporate our framework with a Bézier-boundary gradient approximation strategy. Leveraging this strategy, during the learning process of the 3D scene, we can enable the gradient to properly backpropagate through the “pre-scissoring” operation. Below, we first describe the (forward) image rendering process of DisC-GS, and then explain the Bézier-boundary gradient approximation strategy.
@@ 4.1 Discontinuity-aware Image Rendering
Note that here, since a column of a rotation matrix is already a unit vector, we can omit the normalization of $r_{1}$ and $r_{2}$ and directly transpose them. (2) After deriving $c^{3D}_{curve}\in\mathbb{R}^{4M\times 3}$ storing the $4M$ points in the 3D space coordinate system, we can project each point in $c^{3D}_{curve}$ onto the image plane similar to what we have done in Eq. 1 as:
@@
At this point, for each Gaussian projected onto the image plane via Eq. 1, we have gotten the control points of its desired cubic-Bézier-curves-represented boundary, stored in $c^{2D}_{curve}\in\mathbb{R}^{4M\times 2}$.
@@
where $g_{i}(\cdot)$ represents the indicator function w.r.t. the $i$-th projected Gaussian. Note that via the above modified $\alpha$-blending function, for Gaussians that no longer overlap with the pixel $p$ due to the “scissoring” operation, we can zero out their contributions during calculating the color of $p$.
@@ 4.2 Bézier-boundary Gradient Approximation Strategy
The big picture of our proposed strategy. To tackle the above problems and thus enable Gaussian Splatting to seamlessly render discontinuities, in our framework, we aim to further keep the “differentiability” of the whole discontinuity-aware rendering process. In other words, w.r.t. the discontinuous indicator function $g(\cdot)$ that is newly incorporated into the rendering process, we aim to approximate its gradients (partial derivatives) over the control point coordinates stored in $c^{2D}_{curve}$, in a way that enables the approximated gradients to effectively guide the learning process of the 3D scene representation. To achieve this, we propose a Bézier-boundary gradient approximation strategy. Below, to ease our explanation of the strategy, we focus on discussing how we approximate $\frac{\partial g}{\partial c^{2D}_{curve}[0,0]}$, i.e., the partial derivative of the indicator function $g(\cdot)$ over the $x$ coordinate of the first control point stored in $c^{2D}_{curve}$. Note that the application of the strategy to the remaining coordinates stored in $c^{2D}_{curve}$ follows a similar process (more details are provided in Supplementary). Specifically, to approximate $\frac{\partial g}{\partial c^{2D}_{curve}[0,0]}$, based on the chain rule and according to Eq. 9, denoting $g_{sc}(c^{2D}_{curve}[0],c^{2D}_{curve}[1],c^{2D}_{curve}[2],c^{2D}_{curve}[3];p)$ to be $g_{sc}^{0}(p)$, we first have:
@@
In this system of equations, since $x_{p}$, $y_{p}$, and the coordinates in $c^{2D}_{curve}$ all have known values, we initially regard the second equation in the system as a cubic equation w.r.t. $t$, as $t$ is now the only unknown variable in this equation. After solving this cubic equation and with $t$ also known, we can then regard the first equation in the system as a cubic equation w.r.t. $\phi$ and solve it. Finally, by solving the above two equations, we obtain $S_{\phi}$ as the set of all possible real number solutions for $\phi$. Based on the solutions’ scenarios within $S_{\phi}$, we approximate $\frac{\partial g_{sc}^{0}(p)}{\partial c^{2D}_{curve}[0,0]}$ in three different ways below.
@@ 5 Experiments
Implementation details. We conduct our experiments on an RTX 3090 GPU. For the newly introduced attribute $c_{curve}\in\mathbb{R}^{4M\times 2}$, we set its initial learning rate to 2e-4, and set the hyperparameter $M$ to 3. More details are in supplementary.
@@ 5.1 Experimental Results
In Tab. 1, we compare our approach with existing novel view synthesis methods evaluated on the same 13 3D scenes and report the PSNR, SSIM, and LPIPS results. Our framework consistently outperforms other methods on all three metrics and across various datasets, showing its effectiveness. We also show qualitative results in both Fig. 1(c) and Supplementary. As shown, whether representing the 3D scene through 3D or 2D Gaussian distributions, the conventional Gaussian Splatting technique often struggles to render boundaries and discontinuities clearly and with high quality. In contrast, our framework can achieve good rendering quality, even in regions of the image containing numerous boundaries and discontinuities. This further underscores the efficacy of our approach.
@@ 5.2 Ablation Studies
Impact of representing the scene with 2D or 3D Gaussians in DisC-GS. In Sec. 4, we focus on introducing how we apply DisC-GS on 2D Gaussian Splatting. In fact, via simply introducing each Gaussian with $c^{3D}_{curve}$ instead of $c_{curve}$ as its additional attribute and storing the $4M$ points directly in the 3D space coordinate system, we can also easily apply our framework on 3D Gaussian Splatting. Here to verify the generality of our framework, we test also applying our framework on 3D Gaussian Splatting. As shown in Tab. 2, our framework, when applied on both 2D and 3D Gaussian Splattings, can consistently achieve performance improvements, demonstrating the generality of our framework.

### v2

@@ 2 Related Work
Novel View Synthesis. Owing to the wide range of applications, the task of novel view synthesis has received lots of research attention [22, 39, 38, 41, 23, 35, 3, 4, 5, 42, 48, 7, 11, 17, 18, 36, 29, 25, 13, 47, 19, 44, 21, 34, 32, 31, 46, 50, 16]. In the early days, with the emergence of CNN, different works have been proposed to leverage CNN in this task from different perspectives. Among them, Hedman et al. [22] proposed to use CNN to predict blending weights, and Sitzmann et al. [39] proposed to seek help from CNN in performing volumetric ray-marching. As time passed, NeRF tends to become a popular way in representing 3D scenes. Specifically, the original version of NeRF is first proposed by Mildenhall et al. in [35] and after it comes out, a variety of different NeRF-based methods have been further proposed, such as Mip-NeRF [3], NeRF++ [48], and Point-NeRF [42]. Despite the increased efforts, a weakness of NeRF-based methods can be that, to render novel-view images in high visual quality, they often still require a slow rendering process [29, 13]. This can negatively affect the usage of these methods in many real-world scenarios.
@@ 3 Preliminary
where $N$ represents the number of projected Gaussians that overlap $p$, $c_{i}$ represents the color of the $i$-th Gaussian calculated from its corresponding SH coefficients, $\alpha_{i}$ represents the opacity of the $i$-th Gaussian, $\mu^{2D}_{i}$ represents the center of the $i$-th projected Gaussian, and $\Sigma^{2D}_{i}$ represents the covariance matrix of the $i$-th projected Gaussian. Note that, no matter whether Gaussian Splatting represents the 3D scene using 3D or 2D Gaussian distributions, the above equations can describe its rendering process consistently. In fact, as also mentioned in [25], the difference between rendering images from 3D or 2D Gaussians can be reduced to that, when the scene is represented through 2D Gaussians, the scale matrix $S$ of each of the 2D Gaussians should contain a zero column vector. In this work, we apply our framework to both 2D and 3D Gaussian Splattings, achieving performance improvements as shown in Tab. 2. Yet, as pointed out by [25], using 3D Gaussians instead of 2D Gaussians to represent the scene can result in a viewpoint inconsistency problem. Thus, in Sec. 4, we first focus on explaining how our framework is applied to 2D Gaussian Splatting, in which we fix the last column of the scale matrix $S$ of all the Gaussians to be a zero vector. We then discuss the application of our framework on 3D Gaussian Splatting in Sec. 4.3.
@@ 4 Proposed Method: DisC-GS
Yet, we here argue that Gaussian Splatting has a key intrinsic limitation: it may fail to render discontinuities and boundaries accurately. To tackle this problem, in this work, inspired by [52, 27], we propose a novel framework named DisC-GS, which can seamlessly equip Gaussian Spatting with the discontinuity rendering ability. Specifically, during rendering images from the 3D scene, to render discontinuities properly, DisC-GS enables each Gaussian distribution projected onto the image plane to be first “pre-scissored” along certain desired boundaries before being rendered. However, such a “pre-scissoring” operation by itself can break the differentiability of the framework. Considering this, we further incorporate our framework with a Bézier-boundary gradient approximation strategy. Leveraging this strategy, during the learning process of the 3D scene, we can enable the gradient to properly backpropagate through the “pre-scissoring” operation. Below, we first describe the (forward) image rendering process of DisC-GS, and then explain the Bézier-boundary gradient approximation strategy.
@@ 4.1 Discontinuity-aware Image Rendering
where $\mu$ is the center of the Gaussian distribution. Note that here, since a column of a rotation matrix is already a unit vector, we can omit the normalization of $r_{1}$ and $r_{2}$ and directly transpose them. (2) After deriving $c^{3D}_{curve}\in\mathbb{R}^{4M\times 3}$ storing the $4M$ points in the 3D space coordinate system, we can project each point in $c^{3D}_{curve}$ onto the image plane similar to what we have done in Eq. 1 as:
@@
where $P$ represents the projective transformation matrix, and $W$ represents the viewing transformation matrix. At this point, for each Gaussian projected onto the image plane via Eq. 1, we have gotten the control points of its desired cubic-Bézier-curves-represented boundary, stored in $c^{2D}_{curve}\in\mathbb{R}^{4M\times 2}$.
@@
where $g_{i}(\cdot)$ represents the indicator function w.r.t. the $i$-th projected Gaussian. Besides, same as in Eq. 2, $N$ represents the number of projected Gaussians that overlap $p$, $c_{i}$ represents the color of the $i$-th Gaussian calculated from its corresponding SH coefficients, $\alpha_{i}$ represents the opacity of the $i$-th Gaussian, $\mu^{2D}_{i}$ represents the center of the $i$-th projected Gaussian, and $\Sigma^{2D}_{i}$ represents the covariance matrix of the $i$-th projected Gaussian. Note that via the above modified $\alpha$-blending function, for Gaussians that no longer overlap with the pixel $p$ due to the “scissoring” operation, we can zero out their contributions during calculating the color of $p$.
@@ 4.2 Bézier-boundary Gradient Approximation Strategy
The big picture of our proposed strategy. To tackle the above problems and thus enable Gaussian Splatting to seamlessly render discontinuities, in our framework, we aim to further keep the “differentiability” of the whole discontinuity-aware rendering process. In other words, w.r.t. the discontinuous indicator function $g(\cdot)$ that is newly incorporated into the rendering process, we aim to approximate its gradients (partial derivatives) over the control point coordinates stored in $c^{2D}_{curve}$, in a way that enables the approximated gradients to effectively guide the learning process of the 3D scene representation. To achieve this, inspired by [28], we propose a Bézier-boundary gradient approximation strategy. Below, to ease our explanation of the strategy, we focus on discussing how we approximate $\frac{\partial g}{\partial c^{2D}_{curve}[0,0]}$, i.e., the partial derivative of the indicator function $g(\cdot)$ over the $x$ coordinate of the first control point stored in $c^{2D}_{curve}$. Note that the application of the strategy to the remaining coordinates stored in $c^{2D}_{curve}$ follows a similar process (more details are provided in Supplementary). Specifically, to approximate $\frac{\partial g}{\partial c^{2D}_{curve}[0,0]}$, based on the chain rule and according to Eq. 9, denoting $g_{sc}(c^{2D}_{curve}[0],c^{2D}_{curve}[1],c^{2D}_{curve}[2],c^{2D}_{curve}[3];p)$ to be $g_{sc}^{0}(p)$, we first have:
@@
In this system of equations, since $x_{p}$, $y_{p}$, and the coordinates in $c^{2D}_{curve}$ all have known values, we initially regard the second equation in the system as a cubic equation w.r.t. $t$, as $t$ is now the only unknown variable in this equation. After solving this cubic equation and with $t$ also known, we can then regard the first equation in the system as a cubic equation w.r.t. $\phi$ and solve it. Finally, by solving the above two equations (both in just $O(1)$ time complexity), we obtain $S_{\phi}$ as the set of all possible real number solutions for $\phi$. Based on the solutions’ scenarios within $S_{\phi}$, we approximate $\frac{\partial g_{sc}^{0}(p)}{\partial c^{2D}_{curve}[0,0]}$ in three different ways below.
@@
4.3 DisC-GS on 3D Gaussian Splatting
Above, we focus on describing how we use 2D Bézier curves in our DisC-GS framework and correspondingly apply DisC-GS on 2D Gaussian Splatting. Here in this subsection, we further describe how we use 3D Bézier curves in our DisC-GS framework and apply DisC-GS on 3D Gaussian Splatting. Specifically, the transition from 2D to 3D Bézier curves in DisC-GS requires only two minimal modifications. (1) Firstly, for each Gaussian representing the 3D scene, the control points of its Bézier curves are stored directly in the 3D spatial coordinate system rather than in a 2D subspace. Note that, this modification can be very simply made. Specifically, for each Gaussian in the 3D space in our DisC-GS framework, we only need to use $c_{curve}^{3D}\in\mathbb{R}^{4M\times 3}$ instead of $c_{curve}\in\mathbb{R}^{4M\times 2}$ to represent the control point coordinates of its 3D Bézier curves. In other words, for each 3D Gaussian, we only need to introduce it with $c_{curve}^{3D}$ instead of $c_{curve}$ as its new attribute. (2) Moreover, since we already directly introduce $c_{curve}^{3D}$ as the new attribute for each 3D Gaussian in our framework, during rendering, we omit Eq. 4 above in Sec. 4.1, which originally is used to acquire $c_{curve}^{3D}$ from $c_{curve}$. Overall, the above two modifications are sufficient to incorporate DisC-GS with 3D instead of 2D Bézier curves.
@@ 5 Experiments
Implementation details. We conduct our experiments on an RTX 3090 GPU and develop our code mainly based on the GitHub repository [2] provided by Kerbl et al [29]. Moreover, we also get inspired by [34, 50, 44] during our code implementation, and make use of the LPIPS loss during our training process. Furthermore, for the newly introduced attribute $c_{curve}\in\mathbb{R}^{4M\times 2}$, we set its initial learning rate to 2e-4, and set the hyperparameter $M$ to 3. Besides, in the densification procedure of our framework, when a Gaussian is cloned/splitted into two new Gaussians, we assign both the new Gaussians with the same attribute $c_{curve}$ as the original one.
@@ 5.1 Experimental Results
In Tab. 1, we compare our approach (applied on 2D Gaussian Splatting) with existing novel view synthesis methods evaluated on the same 13 3D scenes and report the PSNR, SSIM, and LPIPS results. Our framework consistently outperforms other methods on all three metrics and across various datasets, showing its effectiveness. We also show qualitative results in both Fig. 1(c) and Supplementary. As shown, whether representing the 3D scene through 3D or 2D Gaussian distributions, the conventional Gaussian Splatting technique often struggles to render boundaries and discontinuities clearly and with high quality. In contrast, our framework can achieve good rendering quality, even in regions of the image containing numerous boundaries and discontinuities. This further underscores the efficacy of our approach.
@@ 5.2 Ablation Studies
Impact of representing the scene with 2D or 3D Gaussians in DisC-GS. In Sec. 4.1 and Sec. 3, we focus on introducing how we apply DisC-GS on 2D Gaussian Splatting. After that, in Sec. 4.3, we introduce how DisC-GS can be applied on 3D Gaussian Splatting in a similar way. Here to verify the generality of our framework, we test applying our framework on both 2D and 3D Gaussian Splatting. As shown in Tab. 2, our framework, when applied on both 2D and 3D Gaussian Splattings, can consistently achieve performance improvements, demonstrating the generality of our framework.
@@ 6 Conclusion
Limitations. While our framework enables Gaussian Splatting to perform discontinuity-aware rendering, we acknowledge that same as existing Gaussian Splatting approaches, our framework still holds certain limitations, such as challenges in rendering large scenes.